Represent the following situations in the form of quadratic equations: (i) The area of a rectangular plot is 528 m2 . The length of the plot (in metres) is one more than twice its breadth. We need to find the length and breadth of the plot.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Let us consider,
Breadth of the rectangular plot = x m
Thus, the length of the plot = (2x + 1) m.
As we know,
Area of rectangle = length × breadth = 528 m^2
Putting the value of length and breadth of the plot in the formula,
we get, (2x + 1) × x = 528
⇒ 2x^2 + x =528
⇒ 2x^2 + x - 528 = 0
Therefore, the length and breadth of plot, satisfies the quadratic equation,
2x2 + x - 528 = 0

Thanks